# Multi-marginal temporal Schrödinger Bridge Matching for video generation from unpaired data

### v1

@@ 3.2.3 Theoretical algorithm
for the backward drift $v_{\phi}$. We summarize in Algorithm 1 our method.
@@ 4 Experiments
For all experiments, we employ Brownian motion $(\sigma_{t}B_{t})_{0\leq t\leq T}$ as the reference measure $\mathbb{Q}$ and T=N1T=N-1 where $N$ is the number of marginals. [^3] More examples for each experiment can be found at mmdsbm.notion.site.
4.1 MMtSBM recovers the known exact OT between Gaussian mixtures
Figure 1: Top row: epoch 0, (only noisy flow matching). Bottom row: epoch $5$, (after MMtSBM training). From left to right: snapshots at times $(0,0.5,1,1.3)$. True marginal times $(t_{0},t_{1},t_{2})=(0,1,2)$. The order of the $3$ true marginals is: $t_{0}=$ dark blue; $t_{1}=$ red; $t_{2}=$ light blue. Generated samples are in green. In the background is the quiver plot of the learned score network.
In this 2D experiment akin to Liu et al. (2022), we used $N=3$ mixtures of two standard Gaussian as marginals. In this configuration the optimal transport between each pair of marginals is known exactly: it is a pure translation of each Gaussian components inside the mixtures. After only the warm-up phase (akin to flow matching (Lipman et al., 2023)), we can see that the learned transport maps mix the Gaussian components of the mixtures, resulting in intersecting trajectories as can be seen in the top row of Figure 1. However, after the SB learning phase of MMtSBM, we can see in the bottom row of Figure 1 that the learned trajectories do not intersect each other anymore and that MMtSBM yields the exact static optimal transport map: pure translations between Gaussians.
@@ 4.2 MMtSBM achieves good usual SB metrics
To quantitatively verify that MMtSBM recovers the correct multi-marginal SB in terms of both 1) static coupling and 2) energy minimization, we extended the now classical ”Moons” and ”8Gaussians” experiments found in Tong et al. (2024) and Shi et al. (2023) to our temporal multi-marginal setting in Table 1. Choosing $N=4$, we considered ($\mathcal{N}$ $\to$ Moons $\to$ $\mathcal{N}$ $\to$ Moons), and ($\mathcal{N}$ $\to$ 8Gaussians $\to$ $\mathcal{N}$ $\to$ 8Gaussians). To assess 1) we report the $\mathcal{W}_{2}$ distance of generations vs test set data at target marginal time(s), averaging along the N1=3N-1=3 target times for MMtSBM and comparing this to the single bridge setting. To assess 2) we report the full path energy $\mathbb{E}\left[\int_{0}^{T}\|v(t,\mathbf{Z}_{t})\|^{2}\,dt\right]$ where $Z_{t}$ is the process simulated along the ODE drift 10.
Setting $\mathcal{W}_{2}$ Path Energy
Moons (single bridge) $0.144{\scriptstyle\pm 0.024}$ $\textit{1.580}{\scriptstyle\pm\textit{0.036}}$
Moons $\times 3$ $4.740$
Moons (MMtSBM) $0.148{\scriptstyle\pm 0.041}$ $5.350{\scriptstyle\pm 0.085}$
8 Gaussians (single bridge) $0.338{\scriptstyle\pm 0.091}$ $\textit{14.810}{\scriptstyle\pm\textit{0.255}}$
8 Gaussians $\times 3$ $44.430$
8 Gaussians (MMtSBM) $0.352{\scriptstyle\pm 0.084}$ $46.920{\scriptstyle\pm 0.285}$
Table 1: Comparison of static coupling ($\mathcal{W}_{2}$ column) and energy minimization (Path Energy column). The rows marked $\times 3$” correspond to the hypothetical case where the energy of a single bridge is simply tripled, and are included as an ideal baseline for comparison with our actual multi-bridge setting.
@@ 4.3 MMtSBM scales to 50d Gaussian transport
We next proceed to scaling our method to dimension $d=50$. We follow the setting of (Shi et al., 2023) and consider a Gaussian-to-Gaussian transport experiment, extended to our multi-marginal case. Specifically, we prescribe four Gaussian marginals at times $t=0,1,2,3$: $\mu_{0}=\mathcal{N}(-0.1\cdot\mathbf{1}_{d},I_{d}),\mu_{1}=\mathcal{N}(0.1\cdot\mathbf{1}_{d},I_{d}),\mu_{2}=\mathcal{N}(-0.1\cdot\mathbf{1}_{d},I_{d}),\mu_{3}=\mathcal{N}(0.1\cdot\mathbf{1}_{d},I_{d})$ where $\mathbf{1}_{d}\in\mathbb{R}^{d}$ denotes the vector of all ones, and $I_{d}$ is the $d\times d$ identity matrix.
Refer to captionRefer to captionRefer to caption
@@
Since no closed-form solution is available for the static multi-marginal SB, we compare our method to the sequence of theoretical results for each pairwise SB (Bunne et al., 2023). As shown in Figure 2, the mean converges rapidly to the prescribed values ($0.1$ or 0.1-0.1) across all four marginals. The variance is slightly more difficult to match: for interior marginals the process tends to overestimate the standard deviation. In contrast, the covariance is consistently well reproduced by our method and remains stable across all three transitions. Interestingly, the covariance converges only after the warmup stage, confirming the added value of the subsequent OT phases. Overall, these results show that MMtSBM scales effectively to the multi-marginal Gaussian setting in $d=50$.
4.4 MMtSBM achieves SOTA results on 100d transcriptomic benchmark
We next evaluate our method on the the TrajectoryNet benchmark (Tong et al., 2020) which uses real single-cell RNA-seq embryoid body differentiation data from Moon et al. (2019). We project RNA counts to their first 100 principal components for each of the $N=5$ marginals. We report in Table 2 the Maximum Mean Discrepency (MMD) and Sliced Wasserstein Distance (SWD).
DMSB MMtSBM (ours)
@@ 4.4 MMtSBM achieves SOTA results on 100d transcriptomic benchmark
Average 0.030 0.160 0.019 0.130
@@
NLSB 0.66 0.54
MIOFlow 0.23 0.35
DMSB 0.03 0.20
MMtSBM (ours) 0.02 0.13
Table 2: Results on test set of embryoid body RNA-seq data ($d=100$). Left table: per-marginal metrics. Right table: average over all target marginals; results from Chen et al. (2023a).
Our method consistently outperforms baselines on all marginals, reducing the average MMD from 0.03 to 0.02 and the SWD from 0.20 to 0.13. This demonstrates that enforcing all marginal constraints simultaneously as we do yields sharper and more consistent interpolations across developmental stages, setting a new state of the art on this benchmark.
@@ 4.5 MMtSBM recovers continuous video dynamics from unpaired data
4.5.1 MNIST experiment
We conducted experiments on the MNIST dataset, transporting digits in decreasing order: $4\to 3\to 2\to 1\to 0$. The algorithm was trained directly in image space, in dimension $28\times 28=784$.
@@ 4.5.1 MNIST experiment
As shown in Figure 3, MMtSBM exhibits clear digit morphing, sometimes reusing pixel structures (e.g., the top of the $3$ to form the top of the $2$), which is what is expected from OT in pixel space. This experiment thus demonstrates that MMtSBM manages to learn a complex temporal OT map in image space directly.
4.5.2 Biotine experiment
The (in-house) ”biotine” dataset consists of 3-channel fluorescence images (GFP, membrane, nucleus) of A549 lung epithelial cells cultured in 384-well plates, treated with biotin, and imaged at 19 discrete time steps. Figure 4 shows the unpaired dynamic we have at hand.
@@ 4.5.2 Biotine experiment
In Figure 5 we can see that MMtSBM successfully recovers a plausible generative coupling between exsting timestamps, preserving both spatial and temporal coherence and yielding smooth, low-noise dynamics that extends the starting test image into a continuous video that respects the global phenotype seen in ground truth data (see Figure 4). To the best of our knowledge, this is the first demonstration of a multi-marginal Schrödinger Bridge performing video generation from purely unpaired data. This provides evidence for both the scalability to very high-dimensional data and for the fidelity to the underlying biological process of MMtSBM.
@@ 5 Discussion
In this work we introduce MMtSBM, a novel method that solves the multi-marginal temporal Schrödinger Bridge problem, adapting Bridge Matching (Shi et al., 2023) to our setting. We demonstrate its theoretical soundness as well as it efficiency. We achieve state-of-the-art results in a common literature benchmark and present for the first time an algorithm that produces temporarily coherent videos from purely unpaired data, hoping to lead to many future applications in the scientific domain.
In future works we would like to investigate other regularizations, such as lifting the process to acceleration space to obtain smoother interpolation trajectories, or exploring other reference processes than the Brownian motion. We also intend to investigate learning the transport map in a VAE latent space, as well as definitively proving our theoretical conjectures. We would also like to explore using the single network theory developed in Bortoli et al. (2024) for efficiency gains.
@@ Appendix B Additional Background
B.1 Other properties on IMFF or MMSB
Proposition B.1 (Markov implies reciprocal).
Any Markov measure on $C([0,T],\mathbb{R}^{d})$ is reciprocal. Hence $P^{\star}\in\mathcal{R}(Q)$. See Proposition 2.3 in Léonard (2012).
Proposition B.2 (Sampling with ODE probability flow).
Given the forward and backward drifts of the multi-marginal Schrödinger bridge, one can simulate trajectories using the probability flow ODE ((Song et al., 2021)):
@@ B.1 Other properties on IMFF or MMSB
Although the score function $\nabla\log p_{t}$ is not directly available, (De Bortoli et al., 2021) show that it can be equivalently recovered by averaging the forward and backward drifts:
@@
B.2 Concrete Algorithms
Algorithm 2 Warmup Phase (our algorithm)
@@ B.2 Concrete Algorithms
2:Output: Warmup parameters $\theta,\phi$
3:Define bridges $\mathcal{B}=\{(t_{i},t_{i+1})\}$, $b\leftarrow B/|\mathcal{B}|$
4:for direction $\in\{\text{forward},\text{backward}\}$ do
5:  for $n\in\llbracket 0,N_{\text{warmup}}\rrbracket$ do
6:   for all $(t_{i},t_{i+1})\in\mathcal{B}$ in parallel do
7:     Sample $(X_{t_{i}},X_{t_{i+1}})\sim(\pi_{t_{i}}\otimes\pi_{t_{i+1}})^{\otimes b}$, $t_{(i)}\sim\text{Unif}[t_{i},t_{i+1}]^{\otimes b}$
8:   end for
9:   Aggregate $X_{\text{init}},X_{\text{final}},t$; Sample $Z\sim\mathcal{N}(0,I)^{\otimes B}$
10:   $X_{t}\leftarrow\mathrm{Interp}_{t}(X_{\text{init}},X_{\text{final}},Z)$ $\triangleright$ cf. equation 11
11:   Update $\theta$ if forward with $\ell^{\text{fwd}}$ equation 13, else $\phi$ with $\ell^{\text{bwd}}$ equation 14
@@
13:end for
Algorithm 3 $M$-IMF Phase (our algorithm)
@@
2:Output: Finetuned parameters $\theta,\phi$
3:Define bridges $\mathcal{B}=\{(t_{i},t_{i+1})\}$, $b\leftarrow B/|\mathcal{B}|$
4:for $N\in\llbracket 0,N_{\text{finetune}}\rrbracket$ do
5:  for all $(t_{i},t_{i+1})\in\mathcal{B}$ in parallel do
6:   Sample $(X_{t_{i}},X_{t_{i+1}})$ from $(\pi_{t_{i}}\otimes\pi_{t_{i+1}})^{\otimes b}$
7:   Sample $t_{(i)}\sim\text{Unif}[t_{i},t_{i+1}]^{\otimes b}$
8:  end for
9:  Aggregate $X_{\text{init}},X_{\text{final}},t$
10:  for direction $\in\{\text{backward},\text{forward}\}$ do
11:   for $n\in\llbracket 0,N_{\text{inner}}\rrbracket$ do
12:     if direction = forward then
13:      $\hat{X}_{\text{init}}\leftarrow\mathrm{SDE}(X_{\text{final}},v_{\phi})$ $\triangleright$ cf. equation 12
14:      $X_{t}\leftarrow\mathrm{Interp}_{t}(\hat{X}_{\text{init}},X_{\text{final}},Z)$ $\triangleright$ cf. equation 11
15:      Update $\theta$ with $\ell^{\text{fwd}}$ equation 13
16:     else
17:      $\hat{X}_{\text{final}}\leftarrow\mathrm{SDE}(X_{\text{init}},v_{\theta})$ $\triangleright$ cf. equation 12
18:      $X_{t}\leftarrow\mathrm{Interp}_{t}(X_{\text{init}},\hat{X}_{\text{final}},Z)$ $\triangleright$ cf. equation 11
19:      Update $\phi$ with $\ell^{\text{bwd}}$ equation 14
20:     end if
@@
23:end for
B.3 Critical Implementation Considerations

### v2

@@ 3.2.3 Theoretical algorithm
for the backward drift $v_{\phi}$.
We summarize in Algorithm 1 our method and provide a practical implementation of IMFF in A.3.
@@ 4 Experiments
For all experiments, we employ Brownian motion $(\sigma_{t}B_{t})_{0\leq t\leq T}$ for the reference measure $\mathbb{Q}$ and T=N1T=N-1 where $N$ is the number of marginals. [^1] Videos for most experiments can be found at mmdsbm.notion.site. All trainings start after a warmup phase like in shi2023diffusionschrodingerbridgematching, detailed in A.3.
4.1 MMtSBM recovers the exact OT between Gaussian mixtures
Figure 1: Top row: epoch 0 (only noisy flow matching). Bottom row: epoch 55 (after MMtSBM training). From left to right: snapshots at times $(0,0.5,1,1.3)$. True marginal times $(t_{0},t_{1},t_{2})=(0,1,2)$. The order of the $3$ true marginals is: $t_{0}=$ dark blue; $t_{1}=$ red; $t_{2}=$ light blue. Generated samples are in green. In the background is the quiver plot of the learned score network.
In this 2D experiment akin to liu2022flowstraightfastlearning, we used $N=3$ mixtures of two standard Gaussian as marginals. In this configuration the optimal transport between each pair of marginals can be computed exactly: it is a pure translation of each Gaussian components inside the mixtures, as we verified with POT (see Figure 6). After only the warm-up phase (akin to flow matching (lipman2023flowmatchinggenerativemodeling)), we can see that the learned transport maps mix the Gaussian components of the mixtures, resulting in intersecting trajectories as can be seen in the top row of Figure 1. However, after the SB learning phase of MMtSBM, we can see in the bottom row that the learned trajectories do not intersect each other anymore and that MMtSBM yields the expected exact optimal transport map: pure translations between Gaussian components.
@@ 4.2 MMtSBM achieves good usual SB metrics
To quantitatively verify that MMtSBM recovers the correct multi-marginal SB in terms of both 1) static coupling and 2) energy minimization, we extended the now classical ”Moons” and ”8Gaussians” experiments found in tong2024improvinggeneralizingflowbasedgenerative and shi2023diffusionschrodingerbridgematching to our temporal multi-marginal setting in Table 1 (see Figure 7). Choosing $N=4$, we considered ($\mathcal{N}$ $\to$ Moons $\to$ $\mathcal{N}$ $\to$ Moons), and ($\mathcal{N}$ $\to$ 8Gaussians $\to$ $\mathcal{N}$ $\to$ 8Gaussians). To assess 1) we report the $\mathcal{W}_{2}$ distance of generations vs test set data at target marginal time(s), averaging along the N1=3N-1=3 target times for MMtSBM and comparing this to the single bridge setting. To assess 2) we report the full path energy $\mathbb{E}\left[\int_{0}^{T}\|v(t,\mathbf{Z}_{t})\|^{2}\,dt\right]$ where $Z_{t}$ is the process simulated along the ODE drift 10.
Model $\mathcal{W}_{2}$ Path Energy
Moons
Single bridge $0.144{\scriptstyle\pm 0.024}$ $\textit{1.580}{\scriptstyle\pm\textit{0.036}}$
Single bridge $\times 3$ $4.740$
MMtSBM (ours) $0.148{\scriptstyle\pm 0.041}$ $5.350{\scriptstyle\pm 0.085}$
8 $\mathcal{N}$
Single bridge $0.338{\scriptstyle\pm 0.091}$ $\textit{14.810}{\scriptstyle\pm\textit{0.255}}$
Single bridge $\times 3$ $44.430$
MMtSBM (ours) $0.352{\scriptstyle\pm 0.084}$ $46.920{\scriptstyle\pm 0.285}$
Table 1: Comparison in terms of static coupling ($\mathcal{W}_{2}$”) and energy minimization (”Path Energy”). The rows marked $\times 3$” correspond to the hypothetical case where the energy of a single bridge is simply tripled, and are included as an ideal baseline for comparison with our actual multi-bridge setting. All metrics apart from ours are from shi2023diffusionschrodingerbridgematching.
@@ 4.3 MMtSBM scales to 50d Gaussian transport
We next proceed to scaling our method to dimension $d=50$. We follow the setting of (shi2023diffusionschrodingerbridgematching) and consider a Gaussian-to-Gaussian transport experiment, extended to our multi-marginal case. Specifically, we prescribe four Gaussian marginals at times $t=0,1,2,3$: $\mu_{0}=\mbox{$\mathcal{N}(-0.1\cdot\mathbf{1}_{d},I_{d})$},\mu_{1}=\mathcal{N}(0.1\cdot\mathbf{1}_{d},I_{d}),\mu_{2}=\mathcal{N}(-0.1\cdot\mathbf{1}_{d},I_{d}),\mu_{3}=\mathcal{N}(0.1\cdot\mathbf{1}_{d},I_{d})$ where $\mathbf{1}_{d}\in\mathbb{R}^{d}$ denotes the vector of all ones, and $I_{d}$ is the $d\times d$ identity matrix.
@@
Since no closed-form solution is available for the static multi-marginal SB, we compare our method to the sequence of theoretical results for each pairwise SB (bunne2022). As shown in Figure 2, the mean converges rapidly to the prescribed values ($0.1$ or 0.1-0.1) across all four marginals. The variance is slightly more difficult to match: for interior marginals the process tends to overestimate the standard deviation. In contrast, the covariance is consistently well reproduced by our method and remains stable across all three transitions. Interestingly, the covariance converges only after the warmup stage, confirming the added value of the subsequent OT phases. Overall, these results show that MMtSBM scales effectively to the multi-marginal Gaussian setting in $d=50$.
4.4 MMtSBM achieves SOTA results on 100d transcriptomic benchmarks
We next evaluate our method on the the Embryoid Body (EB) (Moon2019) and MULTI (pmlr-v176-lance22a) benchmarks, two trajectory inference tasks on real single-cell RNA-seq data. We project RNA counts to their first $d=100$ principal components for each of the $N=5$ and $N=4$ marginals, respectively. [^2] We actually reuse preprocessed data from tong2020trajectorynetdynamicoptimaltransport and tong2024simulationfreeschrodingerbridgesscore. We report the Maximum Mean Discrepency (MMD) and Sliced Wasserstein Distance (SWD) for EB in Table 2, and the Wasserstein-$1$ distance for MULTI in Table 3. For the EB benchmark we train on all marginals, while for the MULTI benchmark we leave-out one of either intermediate times ($t=1$ or $t=2$) during training.
DMSB (chen2023deepmomentummultimarginalschrodinger) MMtSBM (ours)
@@ 4.4 MMtSBM achieves SOTA results on 100d transcriptomic benchmarks
Average 0.032 $\scriptstyle\pm 3e-3$} 0.160 $\scriptstyle\pm 2e-2$} 0.019 $\scriptstyle\pm 4e-4$} 0.130 $\scriptstyle\pm 2e-3$}
@@
NLSB (koshizuka2023neurallagrangianschrodingerbridge) 0.66 0.54
MIOFlow (huguet2022manifoldinterpolatingoptimaltransportflows) 0.23 0.35
DMSB (chen2023deepmomentummultimarginalschrodinger) 0.032 $\scriptstyle\pm 3e-3$} 0.16 $\scriptstyle\pm 2e-2$}
MMtSBM (ours) 0.019 $\scriptstyle\pm 4e-4$} 0.130 $\scriptstyle\pm 2e-3$}
Table 2: MMD and SWD of generations vs test set for the $d=100$ EB benchmark. Our generations start from $\mu_{t=0}^{\text{test}}$. Top table: per-marginal metrics. Bottom table: average over all marginals. Others’ results are from chen2023deepmomentummultimarginalschrodinger. Our error margins are over 10 evaluations while DMSB’s are over 3. Best value in bold.
Method $\mathcal{W}_{1}$ ($\downarrow$)
Schrödinger Bridge
WLF-SB (neklyudov2024computationalframeworksolvingwasserstein) $55.065{\scriptstyle\,\pm 5.499}$
[SF]2M-Exact (tong2024simulationfreeschrodingerbridgesscore) $52.888{\scriptstyle\,\pm 1.986}$
[SF]2M-Geo (tong2024simulationfreeschrodingerbridgesscore) $52.203{\scriptstyle\,\pm 1.957}$
MMtSBM (ours) $\underline{44.542{\scriptstyle\,\pm 0.637}}$
No precomputed OT conditioning
I-CFM (tong2024improvinggeneralizingflowbasedgenerative) $57.262{\scriptstyle\,\pm 3.855}$
[…]
MMtSBM (ours) $\underline{44.542{\scriptstyle\,\pm 0.637}}$
Method $\mathcal{W}_{1}$ ($\downarrow$)
Wasserstein Gradient Flows
WLF-SB (neklyudov2024computationalframeworksolvingwasserstein) $55.065{\scriptstyle\,\pm 5.499}$
WLF-OT (neklyudov2024computationalframeworksolvingwasserstein) $55.416{\scriptstyle\,\pm 6.097}$
WLF-UOT (neklyudov2024computationalframeworksolvingwasserstein) $54.222{\scriptstyle\,\pm 5.827}$
WLF-(OT+potential) (neklyudov2024computationalframeworksolvingwasserstein) $47.365{\scriptstyle\,\pm 0.051}$
WLF-(UOT+potential) (neklyudov2024computationalframeworksolvingwasserstein) $45.231{\scriptstyle\,\pm 0.010}$
Flow Matching with exact OT conditioning
OT-CFM (tong2024improvinggeneralizingflowbasedgenerative) $54.814{\scriptstyle\,\pm 5.858}$
[…]
Metric-aware interpolation with exact OT conditioning
GAGA (sun2025geometryawaregenerativeautoencoderswarped) $\mathbf{27.04{\scriptstyle\,\pm 2.95}}$
Table 3: $\mathcal{W}_{1}$ of generations vs left-out test set for the $d=100$ MULTI benchmark. Generations start from $\mu_{i-1}^{\text{test}}$ where $i$ is the left-out time. Reported figures are the average between left-out $t=1$ and $t=2$ marginals. Our error margin is over 3 training runs. Best value in bold, second best underlined. See A.5.5 for details & comments.
[…]
@@ 4.5 MMtSBM recovers continuous video dynamics from unpaired data
4.5.1 MNIST digit morphing experiment
We conducted experiments on the MNIST dataset of hand-written digits, transporting digits in decreasing order: $4\to 3\to 2\to 1\to 0$. The algorithm was trained directly in image space, in dimension $28\times 28=784$. As shown in Figure 3, MMtSBM exhibits clear digit morphing, sometimes reusing pixel structures (e.g., the top of the $3$ to form the top of the $2$), which is what is expected from OT in pixel space. This experiment thus demonstrates that MMtSBM manages to learn a complex temporal OT map in image space directly.
@@ 4.5.1 MNIST digit morphing experiment
4.5.2 Biotine cell culture experiment
The (in-house) ”biotine” dataset consists of 3-channel fluorescence images (GFP, membrane, nucleus) of A549 lung epithelial cells cultured in 384-well plates, treated with biotin, and imaged at 7 discrete time steps.
Figure 4 shows the unpaired dynamic we have at hand. We can clearly observe fluorescence loss in the cytoplasmic area, corresponding to the green channel. Interestingly, contrary to the above MNIST experiment, a mostly static positional evolution is observed here.
@@ 4.5.2 Biotine cell culture experiment
Time KID ($\downarrow$)
$t=0$ $11.1{\scriptstyle\pm 0.23}$
$t=1$ $13.0{\scriptstyle\pm 0.20}$
$t=2$ $20.1{\scriptstyle\pm 0.23}$
$t=3$ $23.5{\scriptstyle\pm 0.25}$
$t=4$ $26.0{\scriptstyle\pm 0.29}$
$t=5$ $27.7{\scriptstyle\pm 0.31}$
all times $17.1{\scriptstyle\pm 0.32}$
Table 4: KIDs for each marginal and all marginals together, using dinov2-vit-b-14.
This actually stems from the fact that cell position is not a statistically varying information on the biotine dataset, and this (non-existing) signal is thus simply not seen by our purely unpaired method, resulting in non-moving cells. MMtSBM rather reconstructs the OT trajectory in pixel space, yielding very close cellular position while still accurately matching the time-varying phenotype (mainly: the fluorescence loss in the cytoplasm).
We report in Table 4 the KID (bińkowski2021demystifyingmmdgans) values of true vs generated samples, obtained with DINOv2 (oquab2024dinov2learningrobustvisual) as a feature extractor as a baseline reference for future works.
To the best of our knowledge, this is the first demonstration of any method performing video generation from purely unpaired data. Together, this provides evidence for both the scalability to very high-dimensional data and for the fidelity to the underlying biological process of MMtSBM.
@@ 5 Discussion
In this work we introduce MMtSBM, a novel method that solves the multi-marginal temporal Schrödinger Bridge problem, adapting Bridge Matching (shi2023diffusionschrodingerbridgematching) to our setting. We demonstrate the theoretical soundness of both our modeling and algorithm. We show that MMtSBM indeed produces transport maps that are close to the true OT plan in toy experiments and verify its correct behavior in low-dim experiments. We achieve state-of-the-art results in $2$ widely reported single-cell transcriptomic benchmarks, and for the first time demonstrate a method producing temporarily coherent videos from purely unpaired data, hoping to lead to many future scientific applications.
In future works we would like to investigate other regularizations, such as lifting the process to acceleration space to obtain smoother interpolation trajectories, or exploring other empirical reference processes than the Brownian motion. We also intend to investigate learning the transport map in a latent space. We would also like to explore using the single network theory developed in debortoli2024schrodingerbridgeflowunpaired for efficiency gains, as well as simulation-free methods.
LLM Usage Disclosure
LLMs have been used in this work for translation and redaction help, for web search of relevant references and existing literature, and for some annex coding tasks like help on visualizations.
@@ A.1 Additional Background
A.2 Other properties on IMFF or MMSB
Proposition A.1 (Markov implies reciprocal).
Any Markov measure on $C([0,T],\mathbb{R}^{d})$ is reciprocal. Hence $P^{\star}\in\mathcal{R}(Q)$. See Proposition 2.3 in leonard2012.
Proposition A.2 (Sampling with ODE probability flow).
Given the forward and backward drifts of the multi-marginal Schrödinger bridge, one can simulate trajectories using the probability flow ODE ((song2021)):
@@ A.2 Other properties on IMFF or MMSB
Although the score function $\nabla\log p_{t}$ is not directly available, (debortoli2021diffusionschrodingerbridgeapplications) show that it can be equivalently recovered by averaging the forward and backward drifts:
@@
A.3 Concrete Algorithms
We always start trainings with a warmup phase, akin to shi2023diffusionschrodingerbridgematching. It allows MMtSBM to start rectifying the trajectories from a non-random state, which could be complicated because the IMFF phase uses the forward/backward network to train the backward/forward one.
Algorithm 2 Warmup (our algorithm)
@@ A.3 Concrete Algorithms
2:Define bridges $\mathcal{B}=\{(t_{i},t_{i+1})\}$, $b\leftarrow B/|\mathcal{B}|$
3:for direction $\in\{\text{forward},\text{backward}\}$ do
4:  for $n\in\llbracket 0,N_{\text{warmup}}\rrbracket$ do
5:   for all $(t_{i},t_{i+1})\in\mathcal{B}$ in parallel do
6:     Sample $(X_{t_{i}},X_{t_{i+1}})\sim(\pi_{t_{i}}\otimes\pi_{t_{i+1}})^{\otimes b}$, $t_{(i)}\sim\text{Unif}[t_{i},t_{i+1}]^{\otimes b}$
7:   end for
8:   Aggregate $X_{\text{init}},X_{\text{final}},t$; Sample $Z\sim\mathcal{N}(0,I)^{\otimes B}$
9:   $X_{t}\leftarrow\mathrm{Interp}_{t}(X_{\text{init}},X_{\text{final}},Z)$ $\triangleright$ cf. equation 11
10:   Update $\theta$ if forward with $\ell^{\text{fwd}}$ equation 13, else $\phi$ with $\ell^{\text{bwd}}$ equation 14
11:  end for
@@
13:Output: Warmup parameters $\theta,\phi$
Algorithm 3 Iterative Markovian Factorized Fitting (IMFF) (our algorithm)
@@
2:Define bridges $\mathcal{B}=\{(t_{i},t_{i+1})\}$, $b\leftarrow B/|\mathcal{B}|$
3:for $N\in\llbracket 0,N_{\text{finetune}}\rrbracket$ do
4:  for all $(t_{i},t_{i+1})\in\mathcal{B}$ in parallel do
5:   Sample $(X_{t_{i}},X_{t_{i+1}})$ from $(\pi_{t_{i}}\otimes\pi_{t_{i+1}})^{\otimes b}$
6:   Sample $t_{(i)}\sim\text{Unif}[t_{i},t_{i+1}]^{\otimes b}$
7:  end for
8:  Aggregate $X_{\text{init}},X_{\text{final}},t$
9:  for direction $\in\{\text{backward},\text{forward}\}$ do
10:   for $n\in\llbracket 0,N_{\text{inner}}\rrbracket$ do
11:     if direction = forward then
12:      $\hat{X}_{\text{init}}\leftarrow\mathrm{SDE}(X_{\text{final}},v_{\phi})$ $\triangleright$ cf. equation 12
13:      $X_{t}\leftarrow\mathrm{Interp}_{t}(\hat{X}_{\text{init}},X_{\text{final}},Z)$ $\triangleright$ cf. equation 11
14:      Update $\theta$ with $\ell^{\text{fwd}}$ equation 13
15:     else
16:      $\hat{X}_{\text{final}}\leftarrow\mathrm{SDE}(X_{\text{init}},v_{\theta})$ $\triangleright$ cf. equation 12
17:      $X_{t}\leftarrow\mathrm{Interp}_{t}(X_{\text{init}},\hat{X}_{\text{final}},Z)$ $\triangleright$ cf. equation 11
18:      Update $\phi$ with $\ell^{\text{bwd}}$ equation 14
19:     end if
20:   end for
@@
23:Output: Finetuned parameters $\theta,\phi$
A.4 Critical Implementation Considerations